Circular No. 7907

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2002 JN_{16} (LINEAR) An apparently asteroidal object reported by LINEAR (discovery ob-

servation given below), designated 2002 JN_{16}, and having an unusual orbit (e.g., MPEC 2002-J42) has been found by C. A. Trujillo, California Institute of Technology, to be cometary (coma extended 10′′–15′′ from the nuclear condensation in p.a. 270^{o}) on images taken on May 17.2 UT as part of a program by M. Brown and himself with the 1.2-m Schmidt telescope at Palomar. At Trujillo’s request, J. Tonry obtained images of 2002 JN_{16} through clouds with the University of Hawaii 2.2-m reflector at Mauna Kea on May 22.35 that show a similar appearance and m_{1} = 17.2.

2002 UT  α_{2000}  δ_{2000}  m_{2}
May 9.27712  15^{h}33^{m}09.^{s}60  −18^{o}25′36′′.8  18.0

The available astrometry and the following orbital elements appear on MPEC 2002-K38.

T = 2002 July 28.343 TT  ω = 39.^{o}768
e = 0.49107  Ω = 230.110 }2000.0
q = 1.79810 AU  i = 11.560
a = 3.53312 AU n^{o} = 0.148411 P = 6.64 years

1ES 1959+650 D. Horns and A. Konopelko, Max-Planck-Institut für Kernphysik, Hei-

delberg, report, on behalf of the HEGRA collaboration, that follow-up observations of 1ES 1959+650 (cf. IAUC 7903), obtained with the HEGRA system of Cherenkov telescopes at La Palma on May 17–21, show the object in outburst with the overall flux above 1 TeV (2.4 × 10^{26} Hz) increasing from (1 ± 0.5) × 10^{−11} photon cm^{−2} s^{−1} to a peak value of (3.5 ± 0.4) × 10^{−11} photon cm^{−2} s^{−1} during May 19.1–19.2 UT. Within the individual nights, a flux increase with a doubling time of < 1 hr was observed. Observations at other wavelengths are encouraged. HEGRA will continue observations on June 2 with daily monitoring.

SUPERNOVA 2000fv IN MCG –04-35-11 Corrigendum. On IAUC 7906, line 4, for 4′′.2 east read 4′′.2 west

2002 May 22  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green